OpenStudy (anonymous):

Check it: Find average value of function f(x,y)=x^2+y^2 over the square region [0,a]x[0,a] where a>0

OpenStudy (anonymous):

i did :\[\int\limits_{0}^{a}\int\limits_{0}^{a}x^2+y^2dxdy\]

OpenStudy (anonymous):

i got a^4/3 +a^4/3

OpenStudy (anonymous):

The integral is correct, but that gives the SUM of (x^2 + y^2) over the area.

OpenStudy (anonymous):

then i did \[\int\limits_{0}^{a}\int\limits_{0}^{a}dxdy\]

OpenStudy (anonymous):

Ah, yes =) You're getting there.

OpenStudy (anonymous):

that gave me a^2

OpenStudy (anonymous):

Right. It's a square with sidelength a. So area is a^2

OpenStudy (anonymous):

so then divided both integrals

OpenStudy (anonymous):

and i got: 2/3(a^2)

OpenStudy (anonymous):

That matches my answer.